OpenStudy (anonymous):

The probability that a train leaves on time is 0.4. The probability that the train arrives on time and leaves on time is 0.28. What is the probability that the train arrives on time, given that it leaves on time? @ganeshie8

OpenStudy (anonymous):

arrive+leaves=0.28 leaves=0.4

OpenStudy (anonymous):

so arrives=0.12

OpenStudy (anonymous):

Ok, so I just add them together?

OpenStudy (anonymous):

ok so use this....... Let A = leaves on time Let B = arrives on time Thus, A n B arrives on time and leaves on time P(arrives on time given leaves on time) = P(B | A) = P(B n A) / P(A) = P(A n B) / P(A) .......Hope it helps...........!!

ganeshie8 (ganeshie8):